solve for zeros of the following quadratics by factoring
n^2+3n-12=6

Answer by nerdybill(7384)

solve for zeros of the following quadratics by factoring
n^2+3n-12=6
n^2+3n-18= 0
(n+6)(n-3) = 0
n = {-6, 3}